# Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study

### v1

@@ 3.3.2.  Fine-tuning
Because detection must perform reliably across all demographic axes, our evaluation framework also explores the effectiveness of data reweighting Kamiran and Calders (2012) to address imbalances. Let $N$ denote the number of samples and $\mathcal{M}_{\phi}$ the model. For a given instance $i$, its labels $Y_{i}$ form a binary vector of length nine, where $Y_{i}^{m}=1$ if the demographic axis $m$ is targeted and 0 otherwise. The weighted loss is defined as:
$\begin{split}\mathcal{L}_{\text{FT}}=-&\frac{1}{9N}\sum_{i=1}^{N}\sum_{m=1}^{9}w_{i}\Big[\alpha_{m}Y_{i}^{m}\log\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))\\ &+(1-Y_{i}^{m})\log\big(1-\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))\big)\Big],\end{split}$
@@
Models. For encoder models, we consider RoBERTa Liu (2019) and DeBERTa He et al. (2020) and for decoder-only models we consider GPT-2 Radford et al. (2019). For each model, we consider various parameter scales where, across models, the parameters range from 125M to 1.5B.
3.4.  Evaluation metrics
Our comprehensive framework uses metrics capturing three dimensions: (i) distinguishing biased vs. unbiased text, (ii) accurate multi-label classification of bias types, and (iii) ensuring parity across demographic axes and multi-targeted texts.
@@ 3.4.  Evaluation metrics
Multi-label bias detection. Alongside macro $F_{1}^{\text{M}}$ and micro $F_{1}^{\mu}$ scores, we report two multi-label measures Sorower (2010):
@@
  • Per-demographic. Following predictive fairness Hardt et al. (2016); Zafar et al. (2017), we compute the maximum absolute error gap, i.e., overall detection disparity across individual demographic axes: $\Delta_{\mathcal{P}}=\underset{m,m^{\prime}}{\text{max}}\big|\mathcal{P}_{m}-\mathcal{P}_{m^{\prime}}\big|$.
  • Multi-demographic. Inspired by Kearns et al. (2018), we measure if the models make systematically more errors in detecting biases that specifically target multiple axes simultaneously (e.g., {GEN,RAC}) relative to biases that target each constituent axis (e.g., only GEN or RAC). $\mathcal{G}^{\{m,m^{\prime}\}}_{\mathcal{P}}=\underset{x\in\{m,m^{\prime}\}}{\operatorname{max}}\big|\mathcal{P}_{\{m,m^{\prime}\}}-\mathcal{P}_{x}\big|$. This measure helps us understand if the FPR or FNR of multi-axis targeted biased instances is markedly higher, indicating potential blind spots for automated bias detection methods.
@@
4.  Evaluating social bias detection
4.1.  Prompting methods
In-context learning substantially improves detection. Across all models, providing few-shot examples consistently boosts binary $F_{1}$, lowers FNR, and increases multi-label metrics (MR, HL, $F_{1}^{\text{M}}$). Gains are significant with as few as five examples, while moving from five to ten shots yields only marginal improvements. However, inference time grows with the number of examples, highlighting the accuracy–efficiency tradeoff in prompting. Beyond the reported results, we also analyzed alternative in-context setups (in the appendix). We found that (i) RAG-based example selection outperforms random sampling, and (ii) alternative embeddings Youdao (2023) yield comparable results.
@@ 4.1.  Prompting methods
4.2.  Fine-tuning methods
@@
Architecture influences performance. Encoder models (RoBERTa, DeBERTa) consistently outperform decoder models (GPT-2), irrespective of scale. GPT-2-XL underperforms on binary and multi-label detection. In contrast, DeBERTa-v2-XL and RoBERTa-large achieve higher detection scores. Inference times also reflect architectural complexity: decoder models remain faster, whereas DeBERTa-v2-XL is particularly slow due to disentangled attention He et al. (2020).
@@
5.  Evaluating detection disparities
@@
5.1.  Per-demographic axis disparity
@@
5.2.  Multi-demographic disparity
@@
Fine-tuned models show persistent gaps. Despite strong per-axis parity, fine-tuned models underperform on multi-axis instances, reflecting gerrymandering Kearns et al. (2018) in their detection performance. For example, RoBERTa-large with reweighting achieves $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FNR}}\approx 0.436$ and $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FNR}}\approx 0.373$, higher than few-shot Llama-3.1-70B and Qwen-2.5-72B. Encoder models outperform GPT-2, and scaling improves parity (e.g., DeBERTa-v2-XL at $\approx 0.28$ vs. DeBERTa-v3-large at 0.39 for FNR regarding {GEN,SO}). As with per-axis results, reweighting reduces FNR gaps but can slightly raise FPR gaps.
@@ 6.  Discussion
Inclusive stakeholder input. Automation enables large-scale auditing while minimizing human exposure to harmful content Steiger et al. (2021). However, bias detection remains a socio-technical challenge: social biases are culturally situated, and the detection pipeline must be guided by participatory processes Markov et al. (2023).
@@
Data coverage. Our evaluation relies mainly on English datasets reflecting global North contexts, which limits generalizability to other languages and regions. Progress will require multilingual and cross-lingual benchmarks Corazza et al. (2020); Huang and Xiong (2023); Neplenbroek et al. (2024), methods adapted for low-resource settings, and attention to language-specific biases that may not align with English-centric taxonomies. Dataset diversity should also ensure sufficient coverage of social biases that may exist in different cultural contexts. Coverage should also extend to more categories, e.g., political beliefs and biases encoded across multiple data modalities.
@@
Methodological opportunities. Our findings highlight both promises and existing limitations of scalable automated bias detection. Future work should further explore advanced prompting (e.g., chain-of-thought Wei et al. (2022)) and efficient fine-tuning (e.g., using adapters). To handle multimodal data, exploring effective and scalable solutions remains an interesting open challenge. Moreover, while our results offer initial guidance, balancing the potential gains from training and deploying larger models and the scalability required for auditing massive corpora warrants deeper exploration. Finally, our study shows that detecting infrequent, subtle, and multi-axis biases remains a crucial underexplored direction for future research.
7.  Bibliographical References
[…]
  • Barikeri et al. (2021) Soumya Barikeri, Anne Lauscher, Ivan Vulić, and Goran Glavaš. 2021. Redditbias: A real-world resource for bias evaluation and debiasing of conversational language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 1941–1955.
  • Bartl et al. (2020) Marion Bartl, Malvina Nissim, and Albert Gatt. 2020. Unmasking contextual stereotypes: Measuring and mitigating bert’s gender bias. In Proceedings of the Second Workshop on Gender Bias in Natural Language Processing.
  • Blodgett et al. (2020) Su Lin Blodgett, Solon Barocas, Hal Daumé III, and Hanna Wallach. 2020. Language (technology) is power: A critical survey of “bias” in NLP. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5454–5476, Online. Association for Computational Linguistics.
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems, 33:1877–1901.
  • Chen et al. (2024a) Jianfa Chen, Emily Shen, Trupti Bavalatti, Xiaowen Lin, Yongkai Wang, Shuming Hu, Harihar Subramanyam, Ksheeraj Sai Vepuri, Ming Jiang, Ji Qi, et al. 2024a. Class-rag: Real-time content moderation with retrieval augmented generation. arXiv preprint arXiv:2410.14881.
  • Chen et al. (2024b) Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian, and Zheng Liu. 2024b. Bge m3-embedding: Multi-lingual, multi-functionality, multi-granularity text embeddings through self-knowledge distillation. arXiv preprint arXiv:2402.03216.
  • Congress (1964) United States Congress. 1964. Title vii of the civil rights act of 1964. Pub. L. No. 88-352, 78 Stat. 241; codified at 42 U.S.C. § 2000e et seq.
  • Corazza et al. (2020) Michele Corazza, Stefano Menini, Elena Cabrio, Sara Tonelli, and Serena Villata. 2020. A multilingual evaluation for online hate speech detection. ACM Transactions on Internet Technology (TOIT), 20(2):1–22.
  • Dodge et al. (2021) Jesse Dodge, Maarten Sap, Ana Marasović, William Agnew, Gabriel Ilharco, Dirk Groeneveld, Margaret Mitchell, and Matt Gardner. 2021. Documenting large webtext corpora: A case study on the colossal clean crawled corpus. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 1286–1305, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.
  • Dubey et al. (2024) Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Amy Yang, Angela Fan, et al. 2024. The llama 3 herd of models. arXiv preprint arXiv:2407.21783.
  • ElSherief et al. (2021) Mai ElSherief, Caleb Ziems, David Muchlinski, Vaishnavi Anupindi, Jordyn Seybolt, Munmun De Choudhury, and Diyi Yang. 2021. Latent hatred: A benchmark for understanding implicit hate speech. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 345–363.
  • EU FRA (2018) EU FRA. 2018. Handbook on European Non-Discrimination Law. Publications Office of the European Union, Luxembourg.
  • European Commission (2025) European Commission. 2025. Third draft of the general-purpose ai code of practice. Draft prepared by independent experts under the coordination of the European AI Office.
  • European Union (2024) European Union. 2024. Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence. [Accessed: 2025-03-27].
  • Felkner et al. (2023) Virginia Felkner, Ho-Chun Herbert Chang, Eugene Jang, and Jonathan May. 2023. Winoqueer: A community-in-the-loop benchmark for anti-lgbtq+ bias in large language models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 9126–9140.
  • Gallegos et al. (2024) Isabel O. Gallegos, Ryan A. Rossi, Joe Barrow, Md Mehrab Tanjim, Sungchul Kim, Franck Dernoncourt, Tong Yu, Ruiyi Zhang, and Nesreen K. Ahmed. 2024. Bias and fairness in large language models: A survey. Computational Linguistics, 50(3):1097–1179.
  • Gehman et al. (2020) Samuel Gehman, Suchin Gururangan, Maarten Sap, Yejin Choi, and Noah A Smith. 2020. Realtoxicityprompts: Evaluating neural toxic degeneration in language models. arXiv preprint arXiv:2009.11462.
  • GLM et al. (2024) Team GLM, Aohan Zeng, Bin Xu, Bowen Wang, Chenhui Zhang, Da Yin, Dan Zhang, Diego Rojas, Guanyu Feng, Hanlin Zhao, et al. 2024. Chatglm: A family of large language models from glm-130b to glm-4 all tools. arXiv preprint arXiv:2406.12793.
  • Hardt et al. (2016) Moritz Hardt, Eric Price, and Nati Srebro. 2016. Equality of opportunity in supervised learning. Advances in neural information processing systems, 29.
  • Hartvigsen et al. (2022) Thomas Hartvigsen, Saadia Gabriel, Hamid Palangi, Maarten Sap, Dipankar Ray, and Ece Kamar. 2022. Toxigen: A large-scale machine-generated dataset for adversarial and implicit hate speech detection. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 3309–3326.
  • He et al. (2020) Pengcheng He, Xiaodong Liu, Jianfeng Gao, and Weizhu Chen. 2020. Deberta: Decoding-enhanced bert with disentangled attention. arXiv preprint arXiv:2006.03654.
  • Huang and Xiong (2023) Yufei Huang and Deyi Xiong. 2023. Cbbq: A chinese bias benchmark dataset curated with human-ai collaboration for large language models. arXiv preprint arXiv:2306.16244.
  • Hube and Fetahu (2018) Christoph Hube and Besnik Fetahu. 2018. Detecting biased statements in wikipedia. In Companion proceedings of the the web conference 2018, pages 1779–1786.
  • Inan et al. (2023) Hakan Inan, Kartikeya Upasani, Jianfeng Chi, Rashi Rungta, Krithika Iyer, Yuning Mao, Michael Tontchev, Qing Hu, Brian Fuller, Davide Testuggine, et al. 2023. Llama guard: Llm-based input-output safeguard for human-ai conversations. arXiv preprint arXiv:2312.06674.
  • Kamiran and Calders (2012) Faisal Kamiran and Toon Calders. 2012. Data preprocessing techniques for classification without discrimination. Knowledge and information systems, 33(1):1–33.
  • Kearns et al. (2018) Michael Kearns, Seth Neel, Aaron Roth, and Zhiwei Steven Wu. 2018. Preventing fairness gerrymandering: Auditing and learning for subgroup fairness. In International conference on machine learning, pages 2564–2572. PMLR.
  • Klose et al. (2025) Alexander Klose, Doris Liebscher, Maria Wersig, and Michael Wrase, editors. 2025. Landesantidiskriminierungsgesetz Berlin. Nomos, Germany.
  • Koo et al. (2024) Ryan Koo, Minhwa Lee, Vipul Raheja, Jong Inn Park, Zae Myung Kim, and Dongyeop Kang. 2024. Benchmarking cognitive biases in large language models as evaluators. In Findings of the Association for Computational Linguistics ACL 2024, pages 517–545.
  • Kreutzer et al. (2022) Julia Kreutzer, Isaac Caswell, Lisa Wang, Ahsan Wahab, Daan Van Esch, Nasanbayar Ulzii-Orshikh, Allahsera Tapo, Nishant Subramani, Artem Sokolov, Claytone Sikasote, et al. 2022. Quality at a glance: An audit of web-crawled multilingual datasets. Transactions of the Association for Computational Linguistics, 10:50–72.
  • Kumar et al. (2024) Deepak Kumar, Yousef Anees AbuHashem, and Zakir Durumeric. 2024. Watch your language: Investigating content moderation with large language models. In Proceedings of the International AAAI Conference on Web and Social Media, volume 18, pages 865–878.
  • Li et al. (2023) Yunqi Li, Lanjing Zhang, and Yongfeng Zhang. 2023. Fairness of chatgpt. arXiv preprint arXiv:2305.18569.
  • Litt (1961) Edgar Litt. 1961. Jewish ethno-religious involvement and political liberalism. Social Forces, 39(4):328–332.
  • Liu (2019) Yinhan Liu. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692, 364.
  • Luccioni and Viviano (2021) Alexandra Luccioni and Joseph Viviano. 2021. What’s in the box? an analysis of undesirable content in the Common Crawl corpus. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 2: Short Papers), pages 182–189, Online. Association for Computational Linguistics.
  • Maheshwari et al. (2024) Gaurav Maheshwari, Dmitry Ivanov, and Kevin El Haddad. 2024. Efficacy of synthetic data as a benchmark. arXiv preprint arXiv:2409.11968.
  • Markov et al. (2023) Todor Markov, Chong Zhang, Sandhini Agarwal, Florentine Eloundou Nekoul, Theodore Lee, Steven Adler, Angela Jiang, and Lilian Weng. 2023. A holistic approach to undesired content detection in the real world. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 37, pages 15009–15018.
  • Mathew et al. (2021) Binny Mathew, Punyajoy Saha, Seid Muhie Yimam, Chris Biemann, Pawan Goyal, and Animesh Mukherjee. 2021. Hatexplain: A benchmark dataset for explainable hate speech detection. In Proceedings of the AAAI conference on artificial intelligence, volume 35, pages 14867–14875.
  • Nadeem et al. (2021) Moin Nadeem, Anna Bethke, and Siva Reddy. 2021. Stereoset: Measuring stereotypical bias in pretrained language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 5356–5371.
  • Nangia et al. (2020) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel Bowman. 2020. Crows-pairs: A challenge dataset for measuring social biases in masked language models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1953–1967.
  • Navigli et al. (2023) Roberto Navigli, Simone Conia, and Björn Ross. 2023. Biases in large language models: origins, inventory, and discussion. ACM Journal of Data and Information Quality, 15(2):1–21.
  • Neplenbroek et al. (2024) Vera Neplenbroek, Arianna Bisazza, and Raquel Fernández. 2024. Mbbq: A dataset for cross-lingual comparison of stereotypes in generative llms. arXiv preprint arXiv:2406.07243.
  • Palla et al. (2025) Konstantina Palla, José Luis Redondo García, Claudia Hauff, Francesco Fabbri, Andreas Damianou, Henrik Lindström, Dan Taber, and Mounia Lalmas. 2025. Policy-as-prompt: Rethinking content moderation in the age of large language models. In Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency, pages 840–854.
  • Parrish et al. (2022) Alicia Parrish, Angelica Chen, Nikita Nangia, Vishakh Padmakumar, Jason Phang, Jana Thompson, Phu Mon Htut, and Samuel Bowman. 2022. Bbq: A hand-built bias benchmark for question answering. In Findings of the Association for Computational Linguistics: ACL 2022.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog, 1(8):9.
  • Raza et al. (2024) Shaina Raza, Muskan Garg, Deepak John Reji, Syed Raza Bashir, and Chen Ding. 2024. Nbias: A natural language processing framework for bias identification in text. Expert Systems with Applications, 237:121542.
  • Rudinger et al. (2018) Rachel Rudinger, Jason Naradowsky, Brian Leonard, and Benjamin Van Durme. 2018. Gender bias in coreference resolution. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), pages 8–14.
  • Sap et al. (2020) Maarten Sap, Saadia Gabriel, Lianhui Qin, Dan Jurafsky, Noah A Smith, and Yejin Choi. 2020. Social bias frames: Reasoning about social and power implications of language. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5477–5490.
  • Sorower (2010) Mohammad S Sorower. 2010. A literature survey on algorithms for multi-label learning. Oregon State University, Corvallis, 18(1):25.
  • Steiger et al. (2021) Miriah Steiger, Timir J Bharucha, Sukrit Venkatagiri, Martin J Riedl, and Matthew Lease. 2021. The psychological well-being of content moderators: the emotional labor of commercial moderation and avenues for improving support. In Proceedings of the 2021 CHI conference on human factors in computing systems, pages 1–14.
  • Sun et al. (2024) Lichao Sun, Yue Huang, Haoran Wang, Siyuan Wu, Qihui Zhang, Chujie Gao, Yixin Huang, Wenhan Lyu, Yixuan Zhang, Xiner Li, et al. 2024. Trustllm: Trustworthiness in large language models. arXiv preprint arXiv:2401.05561, 3.
  • Tabassi (2023) Elham Tabassi. 2023. Artificial intelligence risk management framework (ai rmf 1.0).
  • Vashney (2022) Kush R Vashney. 2022. Trustworthy machine learning. Independently published.
  • Viprey (2002) Mouna Viprey. 2002. New anti-discrimination law adopted. Eurofound. Published: 3 January 2002.
  • Wang et al. (2023) Boxin Wang, Weixin Chen, Hengzhi Pei, Chulin Xie, Mintong Kang, Chenhui Zhang, Chejian Xu, Zidi Xiong, Ritik Dutta, Rylan Schaeffer, et al. 2023. Decodingtrust: A comprehensive assessment of trustworthiness in gpt models. In NeurIPS.
  • Wang et al. (2024) Song Wang, Peng Wang, Tong Zhou, Yushun Dong, Zhen Tan, and Jundong Li. 2024. Ceb: Compositional evaluation benchmark for fairness in large language models. arXiv preprint arXiv:2407.02408.
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems, 35:24824–24837.
  • Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander M. Rush. 2020. Transformers: State-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 38–45, Online. Association for Computational Linguistics.
  • Yang et al. (2024) An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, et al. 2024. Qwen2. 5 technical report. arXiv preprint arXiv:2412.15115.
  • Youdao (2023) NetEase Youdao. 2023. Bcembedding: Bilingual and crosslingual embedding for rag. https://github.com/netease-youdao/BCEmbedding.
  • Zafar et al. (2017) Muhammad Bilal Zafar, Isabel Valera, Manuel Gomez Rodriguez, and Krishna P Gummadi. 2017. Fairness beyond disparate treatment & disparate impact: Learning classification without disparate mistreatment. In Proceedings of the 26th international conference on world wide web, pages 1171–1180.
  • Zeng et al. (2024) Wenjun Zeng, Yuchi Liu, Ryan Mullins, Ludovic Peran, Joe Fernandez, Hamza Harkous, Karthik Narasimhan, Drew Proud, Piyush Kumar, Bhaktipriya Radharapu, et al. 2024. Shieldgemma: Generative ai content moderation based on gemma. arXiv preprint arXiv:2407.21772.
  • Zhan et al. (2025) Xianyang Zhan, Agam Goyal, Yilun Chen, Eshwar Chandrasekharan, and Koustuv Saha. 2025. Slm-mod: Small language models surpass llms at content moderation. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 8774–8790.
  • Zhao et al. (2018) Jieyu Zhao, Tianlu Wang, Mark Yatskar, Vicente Ordonez, and Kai-Wei Chang. 2018. Gender bias in coreference resolution: Evaluation and debiasing methods. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), pages 15–20.

### v2

@@ 3.3.2 Fine-tuning
Because detection must perform reliably across all demographic axes, our evaluation framework also explores the effectiveness of data reweighting (Kamiran and Calders, 2012) to address imbalances. Let $N$ denote the number of samples and $\mathcal{M}_{\phi}$ the model. For a given instance $i$, its labels $Y_{i}$ form a binary vector of length nine, where $Y_{i}^{m}=1$ if the demographic axis $m$ is targeted and 0 otherwise. The weighted loss is defined as:
$\begin{split}\mathcal{L}_{\text{FT}}=-\frac{1}{9N}\sum_{i=1}^{N}\sum_{m=1}^{9}w_{i}\Big[\alpha_{m}Y_{i}^{m}\log\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))+(1-Y_{i}^{m})\log\big(1-\sigma_{m}(\mathcal{M}_{\phi}(d_{i}))\big)\Big],\end{split}$
@@
Models. For encoder models, we consider RoBERTa (Liu, 2019) and DeBERTa (He et al., 2020) and for decoder-only models we consider GPT-2 (Radford et al., 2019). For each model, we consider various parameter scales where, across models, the parameters range from 125M to 1.5B.
3.4 Evaluation metrics
Our comprehensive framework uses metrics capturing three dimensions: (i) distinguishing biased vs. unbiased text, (ii) accurate multi-label classification of bias types, and (iii) ensuring parity in detection performance across demographic axes and for multi-targeted texts compared to single-targeted texts.
@@ 3.4 Evaluation metrics
Multi-label bias detection. Alongside macro $F_{1}^{\text{M}}$ and micro $F_{1}^{\mu}$ scores, we report two multi-label measures (Sorower, 2010):
@@
  • Per-demographic. Following predictive fairness (Hardt et al., 2016; Zafar et al., 2017), we compute the maximum absolute error gap, i.e., overall detection disparity across individual demographic axes: $\Delta_{\mathcal{P}}=\underset{m,m^{\prime}}{\text{max}}\big|\mathcal{P}_{m}-\mathcal{P}_{m^{\prime}}\big|$.
  • Multi-demographic. Inspired by (Kearns et al., 2018), we measure if the models make systematically more errors in detecting biases that specifically target multiple axes simultaneously (e.g., {GEN,RAC}) relative to biases that target each constituent axis (e.g., only GEN or RAC). $\mathcal{G}^{\{m,m^{\prime}\}}_{\mathcal{P}}=\underset{x\in\{m,m^{\prime}\}}{\operatorname{max}}\big|\mathcal{P}_{\{m,m^{\prime}\}}-\mathcal{P}_{x}\big|$. This measure helps us understand if the FPR or FNR of multi-axis targeted biased instances is markedly higher, indicating potential blind spots for automated bias detection methods.
@@
4 Evaluating social bias detection
4.1 Prompting methods
In-context learning substantially improves detection. Across all models, providing few-shot examples consistently boosts binary $F_{1}$, lowers FNR, and increases multi-label metrics (MR, HL, $F_{1}^{\text{M}}$). Gains are significant with as few as five examples, while moving from five to ten shots yields only marginal improvements. However, inference time grows with the number of examples, highlighting the accuracy–efficiency tradeoff in prompting. Beyond the reported results, we also analyzed alternative in-context setups (in the appendix). We found that (i) RAG-based example selection outperforms random sampling, and (ii) alternative embeddings (Youdao, 2023) yield comparable results.
@@ 4.1 Prompting methods
4.2 Fine-tuning methods
@@
Architecture influences performance. Encoder models (RoBERTa, DeBERTa) consistently outperform decoder models (GPT-2), irrespective of scale. GPT-2-XL underperforms on binary and multi-label detection. In contrast, DeBERTa-v2-XL and RoBERTa-large achieve higher detection scores. Inference times also reflect architectural complexity: decoder models remain faster, whereas DeBERTa-v2-XL is particularly slow due to disentangled attention (He et al., 2020).
@@
5 Evaluating detection disparities
@@
5.1 Per-demographic axis disparity
@@
5.2 Multi-demographic disparity
@@
Fine-tuned models show persistent gaps. Despite strong per-axis parity, fine-tuned models underperform on multi-axis instances, reflecting gerrymandering (Kearns et al., 2018) in their detection performance. For example, RoBERTa-large with reweighting achieves $\mathcal{G}^{\{\text{GEN,SO}\}}_{\text{FNR}}\approx 0.436$ and $\mathcal{G}^{\{\text{GEN,RAC}\}}_{\text{FNR}}\approx 0.373$, higher than few-shot Llama-3.1-70B and Qwen-2.5-72B. Encoder models outperform GPT-2, and scaling improves parity (e.g., DeBERTa-v2-XL at $\approx 0.28$ vs. DeBERTa-v3-large at 0.39 for FNR regarding {GEN,SO}). As with per-axis results, reweighting reduces FNR gaps but can slightly raise FPR gaps.
@@ 6 Discussion
Inclusive stakeholder input. Automation enables large-scale auditing while minimizing human exposure to harmful content (Steiger et al., 2021). However, bias detection remains a socio-technical challenge: social biases are culturally situated, and the detection pipeline must be guided by participatory processes (Markov et al., 2023).
@@
Data coverage. Our evaluation relies mainly on English datasets reflecting global North contexts, which limits generalizability to other languages and regions. Progress will require multilingual and cross-lingual benchmarks (Corazza et al., 2020; Huang and Xiong, 2023; Neplenbroek et al., 2024), methods adapted for low-resource settings, and attention to language-specific biases that may not align with English-centric taxonomies. Dataset diversity should also ensure sufficient coverage of social biases that may exist in different cultural contexts. Coverage should also extend to more categories, e.g., political beliefs and biases encoded across multiple data modalities.
@@
Methodological opportunities. Our findings highlight both promises and existing limitations of scalable automated bias detection. Future work should further explore advanced prompting (e.g., chain-of-thought (Wei et al., 2022)) and efficient fine-tuning (e.g., using adapters). To handle multimodal data, exploring effective and scalable solutions remains an interesting open challenge. Moreover, while our results offer initial guidance, balancing the potential gains from training and deploying larger models and the scalability required for auditing massive corpora warrants deeper exploration. Finally, our study shows that detecting infrequent, subtle, and multi-axis biases remains a crucial underexplored direction for future research.
Mitigation through data bias detection. Bias detection enables more than regulatory reporting; it allows for leveraging the flagged data instances for more responsible model development through targeted data augmentation, filtering, or fine-tuning (Gallegos et al., 2024). However, any remedial action must be guided by rigorous ethical considerations and incorporate human oversight.